A mixture of X & Y contains 60% by weight of X .calculate mole fraction of X & Y in the above mixture . ( molecular weight of X = 60 & Y= 78) plz solve this problem & help me

Grade:12th pass

1 Answers

Rahul kurup
6 years ago
let us take 100 g of mixture.in tis 60%,i.e,60 g will be x.remaining 40 g is y. No.of moles of X=60g/60g(given mass/molar mass)=1 mole .similarly  No.of moles of Y =40g/78g=0.51moles(approx.)
Now x1=n1/(n1+n2)=1/1.51=0.662
x2=1-x1=1-0.662=0.338